# Reasoning Grasping via Multimodal Large Language Model

### v2

@@ IV-C Training Strategy
V Reasoning Grasping Dataset
To train the model for reasoning grasping tasks, we introduce our reasoning grasping dataset, which builds upon the GraspNet-1 billion dataset [13]. GraspNet-1 billion consists of 190 indoor tabletop RGB-D scenes, 88 objects, and 97280 images. Its wide variety of scenes, objects, and grasp poses provides an excellent foundation. We extend the GraspNet-1B dataset with reasoning instructions as well as object parts grasping annotations. We provide some examples in our dataset in Fig. 2.
TABLE I: Comparison on different datasets
@@ V Reasoning Grasping Dataset
LISA[22] RGB N/A N/A N/A 1218
DectGPT [30] RGB N/A N/A N/A 5000
Cornell [19] Rect. RGB-D 240 33 8019 1035
Jacquard [8] Rect. RGB-D 11619 $\sim$ 20 1.1M 54,485
@@
OCID-Grasp [3] Rect. RGB-D 89 $\sim$ 7 75K 1763
@@
V-A Reasoning Instruction Generation
As illustrated in Fig. 3, we developed a unique and detailed method to create reasoning text instructions for our dataset by creating detailed object descriptions, automated generation with GPT-4, and manual review. The generated instructions enable robots to grasp based on nuanced, context-rich queries.
V-A1 Description Creation
The initial step in our instruction generation process involved creating detailed descriptions for each object and part, providing essential information such as functionalities, physical attributes, and typical uses. For instance, a pair of scissors was detailed as “A handheld cutting instrument with two crossing metal blades pivoted together, typically used for cutting paper or fabric. This particular pair has a black handle with yellow inner grips.” These descriptions served as the foundation for generating indirect questions and instructions.
V-A2 Automated Generation with GPT-4
The next step in instruction generation involved utilizing a customized variant of the GPT-4 model. This model was tasked with generating indirect questions and instructions for various objects and their parts. The output consists of 10 strings: 5 indirect questions that reference the object’s functions or design, and 5 indirect instructions that describe actions involving the object or its usage in certain tasks. These are crafted to be specific, relevant, and direct for grasping tasks.
V-A3 Manual Review and Refinement
Post-generation, each set of GPT-generated instructions underwent a manual review and refinement process. The manual intervention allowed us to fine-tune the language, enhancing the quality and applicability of the instructions for the targeted objects and their parts. This semi-automated approach balanced the efficiency of machine-generated content and judgment of human experts.
V-B Part Annotation
V-B1 Part Segmentation
We manually segmented the part point clouds of each object using a specially developed annotation tool. The segmentation process was guided by an understanding of how humans typically interact with and use various parts of an object. For instance, a knife was divided into its blade and handle, reflecting the distinct functional roles of each part.
V-B2 Grasping Pose Annotation
Our dataset’s annotation process includes two key components: pose assignment and pose projection, for both 3D and 2D grasping poses, as shown in Fig. 4.
Pose assignment involves aligning the grasping poses with the nearest part of the object. For each object’s segmented parts, we determine the grasping center of each point. These grasping poses are then assigned to the nearest part, ensuring an accurate association with the specific object part they correspond to.
Pose projection transforms the grasping poses of parts into both 3D and 2D representations. This is achieved using camera matrices and object transformations. In the case of 3D, the original grasping poses are preserved in the spatial domain. For 2D projection, the 3D points of each segmented part are projected onto 2D images. This results in 2D masks that depict the spatial distribution of each part on the image plane. Grasping poses within these masks are then mapped to the corresponding parts in 2D.
V-C Dataset Statistics
In ensuring data quality, low-quality grasping poses were eliminated based on confidence evaluations and objects lacking semantic information were removed. In total, our enhanced dataset comprises an extensive collection of 1,730 instruction pairs, 64 objects, 109 segmented parts, and around 100 million associated grasping poses. Table I shows the statistics comparison of the previous grasping and visual reasoning datasets with ours. Unlike other grasping datasets, we offer detailed reasoning instructions and part-specific grasping information.
VI Experiment
VI-A Implementation Details
VI-B Baselines
VI-B1 CLIP $+$ GR-ConvNet Baseline
In this baseline, we combine a CLIP text encoder with a CNN-based grasp detection model, GR-ConvNet [21]. We first utilize the CLIP text encoder to extract the textual feature from input instructions, and this feature is then integrated into the hidden layers of GR-ConvNet. While CLIP can effectively extract language features relevant to visual content, it may fall short in capturing implicit reasoning compared to the VLM.
VI-B2 LLaVa $\rightarrow$ GR-ConvNet Baseline
This is a modular baseline that integrates the latest LLaVA-1.6-34b [25] with a pre-trained grasping detection model GR-ConvNet. As LLaVA cannot directly output the grasping pose, this baseline operates in two stages to obtain the grasping pose: first, LLaVA takes an image and an instruction as the input and is prompted to output a bounding box of the target object in the image. Then GR-ConvNet detects the optimal grasping pose within this specified bounding box. Note that LLaVA-1.6-34b has a much higher performance than LLaVa-v0-7b [24], which is the base model of our proposed reasoning grasping model. This baseline mirrors a similar process to LAN-grasp.
VI-C Results
VI-C1 Text Generation
We first evaluate the precision of text generation by the fine-tuned multi-modal LLM in the reasoning module. Accurately generating special tokens and target names is crucial, since this identified grasping target will be utilized in the subsequent grasping module. The special tokens and target names act as intermediate states and enable the model to eventually generate a precise grasping pose from given instructions. We differentiate between explicit instructions, which directly mention the object’s name, and implicit instructions, which hint at the target object without stating its name. Note that the instructions used for testing were not seen during the training phase. Results, presented in Table II, show that our model can successfully interpret $92.52\%$ and $88.79\%$ of implicit instructions for object grasping and part grasping, respectively. This demonstrates our model’s good reasoning capabilities in understanding the grasping instructions.
TABLE II: Text generation accuracy.
Explicit Instruction Implicit Instruction
Object Grasping 99.01% 92.52%
Part Grasping 95.33% 88.79%
TABLE III: Grasp prediction accuracy on the reasoning grasping dataset.
Object Grasping Object Grasping Part Grasping Part Grasping
Models R@1 R@2 R@3 R@1 R@2 R@3 R@1 R@2 R@3 R@1 R@2 R@3
LLaVA $\rightarrow$ GR-ConvNet 26.80% 39.17% 44.66% 18.34% 23.76% 31.29% 15.38% 23.52% 37.93% 0.92% 3.84% 8.53%
CLIP $+$ GR-ConvNet 50.40% 62.88% 66.38% 44.35% 57.15% 61.09% 43.78% 55.71% 59.61% 28.88% 42.34% 47.37%
Reasoning Grasping 64.49% 84.11% 86.92% 63.55% 73.83% 77.57% 59.81% 70.01% 81.31% 61.68% 71.96% 83.18%
LLaVA $\rightarrow$ GR-ConvNet (with depth) 31.88% 40.72% 45.95% 20.39% 25.77% 34.84% 20.11% 29.30% 36.63% 1.36% 2.77% 11.95%
CLIP $+$ GR-ConvNet (with depth) 54.11% 66.50% 69.57% 45.89% 58.11% 62.50% 45.53% 59.51% 62.27% 33.06% 48.55% 52.25%
Reasoning Grasping (with depth) 60.75% 71.95% 76.63% 57.94% 70.09% 77.46% 69.16% 72.90% 77.55% 64.48% 75.70% 80.37%
VI-C2 Reasoning Grasping Result
Table III shows the grasp prediction results on the reasoning grasping dataset. Performance is reported using the rectangle evaluation metric [18]. A grasp pose is deemed valid if it fulfills the following two conditions: 1) The Intersection over Union (IoU) score between the predicted and ground truth rectangles exceeds $25\%$; 2) The angular deviation between the orientations of the predicted and ground truth rectangles is less than 30 degrees. We reported R@$k$, indicating the presence of valid grasps within the top-$k$ grasp predictions. Four distinct scenarios are evaluated, categorized by the combination of instruction types (explicit or implicit) and grasping targets (object or part). We assess both baseline models and our reasoning grasping model with and without depth input. Note that when depth information is utilized in the reasoning grasping model, it only serves as input for the grasping module and is not incorporated into the reasoning module, as shown in Fig. 1.
VI-C3 Visual Reasoning Ability
Besides reasoning grasping, we also examine the visual reasoning capabilities of our model. The objective is to determine whether the model, after being fine-tuned, maintains the visual comprehension skills of the original LLaVA model. For a quantitative evaluation of performance, we employ GPT-4 to assess the quality of responses generated by our model. This is inspired by the evaluation methods in [24, 7]. For each given image and question pair, both the original LLaVa model and our fine-tuned version provide answers related to the input image. We subsequently submit the questions, and answers from both LLaVA and our model, along with the ground-truth text descriptions, to a text-only GPT-4-based judge. This judge assesses the responses based on their helpfulness, relevance, accuracy, and level of detail in comparison to the ground-truth descriptions. GPT-4 then allocates an overall performance score ranging from 0 to 10, where higher scores denote superior performance.
The image-question-answer sets used in this evaluation are randomly chosen from the LLaVa-Instruct-150k dataset [24]. According to the judge of GPT-4, the original LLaVA-7B-v0 achieves $8.25/10$ when taking ground-truth textual descriptions as references. On the other hand, our model achieves $7.43/10$ under the same GPT-4 judge. This shows that our fine-tuned model still preserves a good portion of the reasoning ability of the original LLaVA. It also shows that the multi-round conversation ability is preserved, as the image-question-answer sets include sequences of multi-round conversations.
VI-D Real World Experiments
Figure 5: Real world experiments with implicit instructions.
We conduct real-world experiments to evaluate the proposed reasoning grasping model and the CLIP $+$ GR-ConvNet baseline. A 7-DoF Franka Emika Panda equipped with a Franka Hand parallel gripper was utilized for grasping execution. To perceive the objects, one Azure Kinect was positioned in front of the robot. A cropped $480\times 480$ RGB image of the workspace was utilized as the input to the model. To execute the 2D grasps generated from the model, we always apply a top-down grasp with the height computed using the depth of the grasping point. In our experiments, we randomly place 4-8 objects on a tabletop and provide either explicit or implicit language instruction to the robot to grasp an object or a specific part, as shown in Fig. 5. We select objects that are either identical or similar to those in the dataset. We deliberately excluded any objects or parts that are not infeasible for grasping, such as items too wide for the robot’s gripper. The performance is evaluated using three evaluation metrics: the correctness of generating special tokens and grasping target names, the accuracy of the output grasp pose, and the success in lifting the object. The results are reported in Table IV.
Overall our proposed model can generate accurate special tokens and target names 39 out of 40 trials. This again shows the excellent reasoning capabilities of our model to interpret the various instructions. The failure case involved the model incorrectly identifying a screwdriver instead of scissors in response to a prompt requesting a tool to open a package box. We classified it as a failure solely because it did not match the specific pairing in our dataset, which serves as the basis for our evaluation metrics. This example underscores that even when the model’s response is technically incorrect, it is still reasonable. In terms of grasping performance, we assessed top-3 output grasp poses and manually selected the most feasible one for execution on the robot arm. We conducted 10 trials for each scenario, totaling 80 trials, to compare reasoning grasping with CLIP $+$ GR-ConvNet baseline. The CLIP $+$ GR-ConvNet baseline and our reasoning grasping model produced 15/40 and 25/40 accurate grasping poses, respectively, with the success rates for lifting objects are 9/40 and 16/40, respectively. The main reasons for accurate poses but fail to lift were due to the object slipping from the gripper or the object being too heavy. The experiment results demonstrate that our model outperforms the baseline in four distinct scenarios, showing its superior ability in reasoning grasping tasks.
TABLE IV: Real-world experiment results.
CLIP + GR-ConvNet Reasoning Grasping
Object Grasping with Explicit Instruction
Token Accuracy
N/A 10/10
Pose Accuracy
5/10 7/10
Execution Success
2/10 5/10
Object Grasping with Implicit Instruction
Token Accuracy
N/A 9/10
Pose Accuracy
3/10 5/10
Execution Success
2/10 3/10
Part Grasping with Explicit Instruction
Token Accuracy
N/A 10/10
Pose Accuracy
5/10 7/10
Execution Success
3/10 4/10
Part Grasping with Implicit Instruction
Token Accuracy
N/A 10/10
Pose Accuracy
2/10 6/10
Execution Success
2/10 4/10
VII Conclusion
Appendix A Object and Part Statistics
TABLE V: Selected Objects and Parts for the Training Split of the Dataset
@@ Appendix A Object and Part Statistics
0 Cracker Box - 10971 4672 35
1 Tomato Soup Can - 16625 8284 34
5 Banana Stem, Flesh 1298 2941 25
7 Red Mug Handle, Rim, Body 3363 2045 25
8 Power Drill Handle, Body, Batterypack, Chuck 2445 4476 26
@@ 11 Strawberry - 130 903 24
14 Peach - 960 1922 25
15 Pear - 948 2069 26
17 Orange - 581 1466 32
18 Knife Handle, Blade 1975 1526 29
20 Red Screwdriver Handle, Shaft 1191 1910 29
@@ 21 Racquetball - 443 573 6
22 Blue Cup Rim, Body 3960 2770 29
36 Daobao Wash Soap Cap, Body 2594 1744 25
37 Nzskincare Mouth Rinse Cap, Body 1806 1786 24
38 Daobao Sod Cap, Body 3256 2227 25
40 Kispa Cleanser Cap, Body 931 1654 25
41 Darlie Toothpaste Cap, Body 442 1181 25
@@ 43 Baoke Marker Cap, Body 953 697 21
44 Hosjam Toothpaste Pump Pump, Body 725 1958 22
@@ 48 Camel Legs, Head, Body 96 663 25
51 Elephant Legs, Head, Body, Truck 218 1692 24
52 Rhinocero Legs, Head, Body, Horn 152 1419 22
57 Weiquan Chicken Bouillon Cap, Body 2297 2617 25
58 Darlie Toothpaste Box - 5526 1646 25
@@ 60 Black Mouse - 170 618 25
61 Dabao Facewash Cap, Body 4683 2297 22
62 Pantene Shampoo Cap, Body 502 1175 26
63 Head Shoulders Supreme Cap, Body 777 1700 24
66 Head Shoulders Care Cap, Body 1338 1269 24
70 Tape - 4465 1859 25
Total 33 54 76377 64355 100
Appendix B 3D Object and Part Visualization
Figure 6: 3D Visualization of Selected Objects and its Parts
Appendix C Prompt for Instruction Generation

### v3

@@ Appendix F Examples of Reasoning Grasping Dataset
Appendix G Comparison with Different Datasets
Table 4: Comparison with different datasets
@@ Appendix G Comparison with Different Datasets
LISA[51] RGB N/A N/A N/A 1,218
DectGPT [52] RGB N/A N/A N/A 5,000
Cornell [23] Rect. RGB-D 240 33 8,019 1,035
Jacquard [24] Rect. RGB-D 11,619 $\sim$ 20 1.1M 54,485
@@
OCID-Grasp [53] Rect. RGB-D 89 $\sim$ 7 75K 1,763
@@
Appendix H Object and Part Statistics
Table 5: Selected Objects and Parts for the Training Split of the Dataset
@@ Appendix H Object and Part Statistics
0 Cracker Box - 10,971 4,672 35
1 Tomato Soup Can - 16,625 8,284 34
5 Banana Stem, Flesh 1,298 2,941 25
7 Red Mug Handle, Rim, Body 3,363 2,045 25
8 Power Drill
Handle, Chuck,
2,445 4,476 26
Battery Pack, Body
@@ 11 Strawberry - 130 903 24
14 Peach - 960 1,922 25
15 Pear - 948 2,069 26
17 Orange - 581 1,466 32
18 Knife Handle, Blade 1,975 1,526 29
20 Red Screwdriver Handle, Shaft 1,191 1,910 29
@@ 21 Racquetball - 443 573 6
22 Blue Cup Rim, Body 3,960 2,770 29
36 Daobao Wash Soap Cap, Body 2,594 1,744 25
37 Nzskincare Mouth Rinse Cap, Body 1,806 1,786 24
38 Daobao Sod Cap, Body 3,256 2,227 25
40 Kispa Cleanser Cap, Body 931 1,654 25
41 Darlie Toothpaste Cap, Body 442 1,181 25
@@ 43 Baoke Marker Cap, Body 953 697 21
44 Hosjam Toothpaste Pump Pump, Body 725 1,958 22
@@ 48 Camel Legs, Head, Body 96 663 25
51 Elephant
Legs, Head,
218 1,692 24
Body, Truck
52 Rhinocero
Legs, Head,
152 1419 22
Body, Horn
57 Weiquan Chicken Bouillon Cap, Body 2,297 2,617 25
58 Darlie Toothpaste Box - 5,526 1,646 25
@@ 60 Black Mouse - 170 618 25
61 Dabao Facewash Cap, Body 4,683 2,297 22
62 Pantene Shampoo Cap, Body 502 1,175 26
63 Head Shoulders Supreme Cap, Body 777 1,700 24
66 Head Shoulders Care Cap, Body 1,338 1,269 24
70 Tape - 4,465 1,859 25
Appendix I Hyperparameters
Appendix J Experiments on Visual Reasoning Ability.
To evaluate the visual reasoning capabilities of our model, we employ a GPT-4 judge to assess the quality of responses generated by our model. For each given image and question pair, both the original LLaVa model and our fine-tuned version provide answers related to the input image. We subsequently submit the questions, and answers from both LLaVA and our model, along with the ground-truth text descriptions, to a text-only GPT-4-based judge. This judge assesses the responses based on their helpfulness, relevance, accuracy, and level of detail in comparison to the ground-truth descriptions. GPT-4 then allocates an overall performance score ranging from 0 to 10, where higher scores denote superior performance. The image-question-answer sets used in this evaluation are randomly chosen from the LLaVa-Instruct-150k dataset [12] with multi-round conversations.